OpenStudy (anonymous):

"Has an x-intercept of 6 and a y-intercept of 3." Could someone write that equation in slope-intercept, point slope, or standard form?

OpenStudy (anonymous):

y = 1/2x + 3

OpenStudy (anonymous):

That's slope-intercept form

OpenStudy (anonymous):

Thank you so much! Could you show me how you got that?

OpenStudy (anonymous):

sure slope-intercept form is y = mx + b m is the slope b is the y-intercept with the x intercept and the y intercept I got a slope of 1/2 then I just inserted it in and got y = 1/2x + 3 since there's a y-intercept of 3

OpenStudy (anonymous):

do you want me to explain slope?

OpenStudy (anonymous):

Sure, I'm not sure how to get the slope.

OpenStudy (anonymous):

ok the slope is change of y / change of x so the x-intercept is (6,0) and the y-intercept is (0,3) so 3-0/0-6 = 3/-6 = 1/-2

OpenStudy (anonymous):

BTW change it to y = 1/-2x + 3 I forgot it in the original answer sorry

OpenStudy (anonymous):

\[y = \frac{ 1 }{-2 }x + 3?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem